OpenStudy (anonymous):

Kelli weighs 414 N, and she is sitting on on a playground swing that hangs 0.5 m above the ground. Her mom pulls the swing back and releases it when the seat is 1.1 m above the ground. If Kelli moves through the lowest point at 2.0 m/s, how much work was done on the swing by friction?

OpenStudy (schrodingers_cat):

Well think of this system in terms of energy conservation. At 1.1m the gravitational potential energy is 414(1.1-.5) = 248.4. All this energy must converted to kinetic energy at the lowest point. So, 248.4 =1/2mv^2 and the difference will be the energy lost to friction. So, m = 414/9.8 = 42.2 kg (.5)(42.2)(4ms)= 84.4 J So, the energy lost to friction is just the difference of 248.4 - 84.4 = 164 J. I hope this helps :)

OpenStudy (anonymous):

Thank you so much!